find three consecutive odd integers such that five times the sum of the first and second is 1 less than 9 times the third
Answer by Earlsdon(6294)

Let the consecutive odd integers be x, x+2 , and x+4, so...
5(x+x+2) = 9(x+4)-1 Simplify.
10x+10 = 9x+36-1 Subtract 9x.
x+10 = 35 Subtract 10.
x = 25, x+2 = 27, and x+4 = 29
The three integers are: 25, 27, and 29.